OpenStudy (anonymous):

What is the sum of the arithmetic sequence 140, 134, 128…, if there are 33 terms?

OpenStudy (anonymous):

difference is -6

OpenStudy (anonymous):

d/2 (a1 + an)

OpenStudy (anonymous):

a1=140 an= 140 -6(32)=

OpenStudy (anonymous):

\[\S_{n} = \ \frac{n(a_{1} + a_{n})}{2} \]

OpenStudy (anonymous):

So you can find a(33) = 140 + (-6)(33 - 1) = 140 - 192 = -52 33(140 - 52)/2 = 1452